Leonard Co is expected to pay a dividend of $2.75. If this dividend is expected to grow at a 3% rate indefinitely and you require a return of 12%,
A. what would you expect to pay for a share of Leonard Co stock?
B. What is the dividend yield for Leonard Co stock?
C. What is the capital gains yield for Leonard Co stock?
Answer #1
A. Share price is the present value of future dividends which is calculated as follows:
Share price = D1/(Ke-g) Where,
= 2.75/(0.12-0.03) D1 = $       2.75
= $    30.56 Ke = 12%
g = 3%
B. Dividend yield = Annual dividend / Current Price
= $       2.75 / $    30.56
= 9.00%
C. Capital Gain Yield = Required return - Dividend yield
= 12% - 9.00%
= 3.00%